The speed of train a is 8 mph slower than the speed of train b. Train a travels 220 miles in the same time it takes train b to travel 260 miles. Find the speed of each train
Answer by checkley77(12844)

D=RT
T=D/R
T=220/(X-8)
T=260/X
220/(X-8)=260/X CROSS MULTIPLY.
260(X-8)=220X
260X-2,080=220X
260X-220X=2,080
40X=2.080
X=2,080/40
X=52 MPH. FOR THE FASTER TRAIN.
52-8=44 MPH. FOR THE SLOWER TRAIN.
pROOF:
220/44=260/52
5=5